Two charges lie on positive x-axis. Charge A (2.0 x 10-9 C) is 4.0 cm from the origin and Charge B is 8.0 cm from the origin. (-3.0 x 10-9 C). What is the total force exerted by these two charges on Charge C (5.0 x 10-9 C) located at the origin?Solve for ForceBonC.



1
Expert's answer
2022-02-25T11:15:51-0500

"F=kq_C(\frac{q_A}{r_A^2}-\frac{q_B}{r_B^2})=35~\mu N."